No idea where to file this under, hopefully someone will be able to help anyway.
x²-6x+9=100 I would appreciate the steps on how to do it, thank you! :)

Answer by ankor@dixie-net.com(22740)

x²- 6x + 9 = 100
Subtract 100 from both sides
x^2 - 6x + 9 - 100 = 0
:
x^2 - 6x - 91 = 0
You solve this using the quadratic formula, but this will factor to
(x-13)(x+7) = 0
two solutions
x = 13
x = -7
:
:
Check in the original equation x=13
13^2- 6(13) + 9 = 100
169 - 78 + 9 = 100
:
check when x = -7
-7^2 - 6(-7) + 9 = 100
49 + 42 + 9 = 100